[Hallc_running] [New Logentry] Hall C RC Update: Monday, 8/27

gaskelld at jlab.org gaskelld at jlab.org
Mon Aug 27 09:05:01 EDT 2018

Logentry Text:
Yesterday there was a lot of effort to optimize the beam to Hall C.  By late swing an optimal solution had not yet been achieved, but the spot at the target was deemed likely good enough to try a quick verision of the SHMS beam steering test.  During the test, minimal beam motion was observed at the first big BPM downstream of the target (IPM3H08) when ramping the dipole, bender, Q1, and Q3. The shift crew ran into trouble ramping Q2 which was resolved by reducing the ramp rate.  Unfortunately, other issues in the machine delayed beam delivery to Hall C until late OWL shift - when beam did come back briefly, it seems that there was little horizontal motion due to Q2, but some apparent vertical motion - this is unexpected, and perhaps an issue with the rotation angle put into the software for the BPM.

The plan for today is for a concentrated effort to understand the optics in the accelerator (throughout the machine and in Hall C in particular).  CW beam may be delivered to Hall C later, but for the purposes of investigating the beam.  Upon completion of beam related investigations, we will need to perform Jay's test plan for calibrating the response of the big BPMS, and then go back to the SHMS steering study.

Since work today will mostly focus on beam delivery issues, 3rd shift workers for DAY and SWING can be on call (provide contact info to shift leader). Depending on progress with the beam today, OWL shift 3rd's may also be place on call.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3591728
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180827/927480aa/attachment.html>

More information about the Hallc_running mailing list